OpenStudy (anonymous):

I know the answer is x = 27 but can't figure out how to get there. The problem is 3x-5+(x+1)/2=/90

jimthompson5910 (jim_thompson5910):

what's over the 90 on the right side?

OpenStudy (anonymous):

nothing

jimthompson5910 (jim_thompson5910):

Is the equation \[\large 3x-5+\frac{x+1}{2}=90\]

OpenStudy (anonymous):

@jim - yes

jimthompson5910 (jim_thompson5910):

\[\large 3x-5+\frac{x+1}{2}=90\] \[\large 6x-10+x+1=180\] \[\large 7x-9=180\] \[\large 7x=180+9\] \[\large 7x=189\] \[\large x=\frac{189}{7}\] \[\large x=27\]

jimthompson5910 (jim_thompson5910):

In the second step, I multiplied EVERY term by the LCD 2 to clear out the fraction.

OpenStudy (anonymous):

Ohhhh.... Thanks so much!